# Learning Generalized Diffusions using an Energetic Variational Approach

### v1

@@ Abstract
Extracting governing physical laws from computational or experimental data is crucial across various fields such as fluid dynamics and plasma physics. Many of those physical laws are dissipative due to fluid viscosity or plasma collisions. For such a dissipative physical system, we propose two distinct methods to learn the corresponding laws of the systems based on their energy-dissipation laws, assuming either continuous data (probability density) or discrete data (particles) are available. Our methods offer several key advantages, including their robustness to corrupted observations, their easy extension to more complex physical systems, and the potential to address higher-dimensional systems. We validate our approach through representative numerical examples and carefully investigate the impacts of data quantity and data property on the model discovery.
@@ 1 Introduction
Constructing surrogate models and discovering physical laws, often represented by nonlinear partial differential equations (PDEs), are two major data-driven approaches that can help us better understand complex natural phenomena. One of the most widely used methods for discovering physical laws is the physics-informed neural network (PINN) [48]. The idea of PINN is to train neural networks using a loss function based on the underlying partial differential equation and noisy observation data. This approach can be traced back to at least the works in [50, 1, 18]. Another powerful approach for extracting governing physical laws from data is a sparse identification of nonlinear dynamical systems (SINDy) [5]. SINDy has gained popularity due to its interpretability and computational efficiency. The SINDy framework is motivated by the pioneer work [4, 53], which uses symbolic regression to recover physical equations from data. Later, a weak-form version of SINDy was developed for learning PDEs [39, 38] and extended to cover mean-field equations [40] and Hamiltonian systems [41]. Koopman operator theory is also used to establish various data-driven analysis for complex dynamics [6, 59, 28]. Nonparametric regression techniques for learning interaction kernels [34, 32, 29, 15, 42, 33, 14] are developed for various equations. Flow maps [10, 12, 31] and kernel flows [60] for learning dynamical systems are introduced. Probabilistic/statistical methods, including Bayesian inferences, maximum likelihood methods, and Wasserstein distances, are introduced to learn stochastic dynamical systems [13, 45, 36]. More recently, in order to maintain the physical properties (e.g., invariant quantities for conserved systems or dissipation rates for dissipative systems) of the original system while learning the system, various structure-preserving learning strategies are developed to learn Hamiltonian systems [27, 7, 8, 37, 3, 16, 19, 23, 11, 30], energy dissipative systems [62, 61, 25], and, more generally, metriplectic systems [20, 21].
@@
However, most of the existing work establish the learning framework based on the corresponding equations, such as (stochastic) ordinary differential equations (SDE/ODE for short) and PDEs. The learned models may not preserve the fundamental physical law of dissipation. Recently, there are growing interesting of learning thermodynamically consistent physical model from variational principles, such as General Equation for Non-Equilibrium Reversible-Irreversible Coupling (GENERIC) formalism [22, 64] and Onsager principle [26, 61]. These variational principles model complex physical processes by accounting for both energy conservation (in reversible processes) and energy dissipation (in irreversible processes). The key idea behind these variational principle-based learning approaches is to parameterize the physical quantities using neural networks while constructing loss functions based on equations derived from these principles.
The goal of this work is to propose a new learning framework based on the energy-dissipation laws of the target physical systems, without relying on the governing equations. Our proposed methods offer several benefits, including robustness to corrupted observations, easy extensions to more general physical systems, and has the potential to handle higher dimensional systems. In this work, we focus on learning potential function and noise intensity in one-dimensional generalized diffusion to illustrate our method and explore its performance under different settings. Extending the approach to higher-dimensional problems and other physical systems is straightforward. For much higher dimensional problems, we leave this as future work.
@@
The rest of the paper is organized as follows. Section 2 provides a brief introduction to the energetic variational approach for generalized diffusions. In Section 3, we propose two methods for learning the governing laws of the systems based on their energy-dissipation laws, using either continuous data (probability density) or discrete data (SDE particles). Section 4 presents several representative examples to validate the performance of our methods. Finally, we conclude with a brief discussion and propose an alternative approach for learning the system based on another type of discrete data (ODE particles) in Section 5.
@@ 2 Formulation
Before proposing the learning framework, we briefly introduce the energetic variational approach (EnVarA for short) [17] for generalized diffusions, which plays an important role in our proposed learning frameworks in the next section.
Motivated by non-equilibrium thermodynamics, particularly the seminal work of Rayleigh [55] and Onsager [43, 44], a complex system can be described by an energy-dissipation law
@@
where $E^{{\rm total}}$ is the sum of the kinetic energy $\mathcal{K}$ and the Helmholtz free energy $\mathcal{F}$, and $\Delta$ is the rate of energy dissipation. Based on the energy-dissipation law (1), EnVarA is a unique, well-defined way to derive the dynamics of the underlying system using the least action principle (LAP) and the maximum dissipation principle (MDP). To be more specific, for the Hamiltonian part of the system, one can employ the LAP, taking variation of the action functional $\mathcal{A}({\bm{x}})=\int_{0}^{T}\left(\mathcal{K}-\mathcal{F}\right){\rm d}t$ with respect to ${\bm{x}}$ (the trajectory in Lagrangian coordinates) [17, 2], to derive the conservative force, i.e., $\delta\mathcal{A}=\int_{0}^{T}\int_{\Omega}({\rm force}_{\text{iner}}-{\rm force}_{\text{conv}})\cdot\delta{\bm{x}}~{}{\rm d}{\bm{x}}{\rm d}t.$ For the dissipation part, one can apply the MDP, taking the variation of the Onsager dissipation functional $\mathcal{D}$ with respect to the “rate” ${\bm{x}}_{t}$, to derive the dissipative force, i.e., $\delta\mathcal{D}=\int_{\Omega}{\rm force}_{\text{diss}}\cdot\delta{\bm{x}}_{t}~{}{\rm d}{\bm{x}}$, where the dissipation functional $\mathcal{D}=\frac{1}{2}\triangle$ in the linear response regime [44]. Subsequently, the force balance condition connects the conservative force and the dissipation force providing the evolution equation of the studied system
$\frac{\delta\mathcal{D}}{\delta{\bm{x}}_{t}}=\frac{\delta\mathcal{A}}{\delta\bm{x}}.$ (2)
The EnVarA has been successfully applied to build various mathematical models in physics, chemical engineering, and biology [57, 58].
@@
where $W$ is a standard $n$-dimensional Brownian motion, the state variable $\bm{X}_{t}$ and the drift coefficient $\bm{a}$ are two $n$-dimensional vectors, the noise intensity $\sigma$ is a scalar function, and $t\in\mathbb{R}^{+}\cup\{0\}$. If the stochastic integral of (3) is interpreted as backward Itô integral, one may obtain the following Fokker-Planck equation (See (c) in Remark 2.1):
$f_{t}+\nabla\cdot(\bm{a}f)=\frac{1}{2}\nabla\cdot[\sigma^{2}\nabla f],$ (4)
@@
When following the fluctuation-dissipation theorem, one restricts the convection coefficient
@@
The Fokker-Planck equations (4) with the condition (5) can be obtained from variation of the following energy dissipative law:
[…]
@@
Here, ${\bm{u}}$ is a certain average velocity of all stochastic trajectories, and $\mathcal{F}[f]$ is the free energy is defined by
@@
One can derive the evolution equation (4) by the general framework of EnVarA [17]. Note that
[…]
$\begin{cases}&\frac{{\rm d}}{{\rm d}t}{\bm{x}}({\bm{X}},t)={\bm{u}}({\bm{x}}({\bm{X}},t),t),\\ &{\bm{x}}({\bm{X}},0)={\bm{X}}.\\ \end{cases}$ (10)
@@
which is the velocity derived from the energy-dissipation law (6). Combining with the continuity equation $f_{t}+\nabla\cdot(\bm{u}f)=0$, one can obtain the Fokker-Planck equation (4) with ${\bm{a}}$ given by (5). We refer the interested reader to [17, 24] and the references therein for more details.
@@
There are at least three ways to express the energy-dissipation law based on different interpretations of the stochastic integral of (3). To be more specific, writing a Taylor expansion of probability distribution function $f(\bm{x},t)$, one may obtain the following PDEs [17]:
@@
(c) $f_{t}+\nabla\cdot(\bm{a}f)=\frac{1}{2}\nabla\cdot[\sigma^{2}\nabla f]$ if using backward Itô integral, yielding PDE with self-adjoint diffusion term.

### v2

@@ Abstract
Extracting governing physical laws from computational or experimental data is crucial across various fields such as fluid dynamics and plasma physics. Many of those physical laws are dissipative due to fluid viscosity or plasma collisions. For such a dissipative physical system, we propose a framework to learn the corresponding laws of the systems based on their energy-dissipation laws, assuming either continuous data (probability density) or discrete data (particles) are available. Our methods offer several key advantages, including their robustness to corrupted/noisy observations, their easy extension to more complex physical systems, and the potential to address higher-dimensional systems. We validate our approaches through representative numerical examples and carefully investigate the impacts of data quantity and data property on model discovery.
keywords:
Energetic variational approach; Data-driven modeling; Fokker-Planck equation; Fluctuation-dissipation theorem.
[…]
35A15, 65M32, 76M21, 76M30
@@ 1 Introduction
Constructing surrogate models that approximate the behavior of physical systems and discovering physical laws, often represented by nonlinear partial differential equations (PDEs), are two major data-driven approaches that can help us better understand complex natural phenomena. As a concrete example, generalized diffusion, a type of mechanical process involving a conserved quantity, can be described by an energy-dissipation law. Generalized diffusion encompasses a wide range of models across various fields, including the Fokker-–Planck equation(FPE) risken1996fokker , the porous medium equation PME , and the Poisson–Nernst–Planck equation PNP ; LiPNP ; LiuPNP . Constructing surrogate models and discovering physical laws for such systems can serve as effective alternatives or complements to costly equation-based state-of-the-art methods, enabling tasks such as prediction, optimal control, and uncertainty quantification.
@@
One of the most widely used methods for discovering physical laws is the physics-informed neural network (PINN) PINN . The idea of PINN is to train neural networks using a loss function based on the underlying partial differential equation and noisy observation data. This approach can be traced back to at least the works in Yannis1 ; Yannis2 ; Yannis3 . Another powerful approach for extracting governing physical laws from data is a sparse identification of nonlinear dynamical systems (SINDy) SINDy . SINDy has gained popularity due to its interpretability and computational efficiency. The SINDy framework is motivated by the pioneer work bongard2007automated ; schmidt2009distilling , which uses symbolic regression to recover physical equations from data. Later, a weak-form version of SINDy was developed for learning PDEs weakSINDy ; weakSINDy4PDE and extended to cover mean-field equations weakSINDy4meanfield and Hamiltonian systems weakSINDy4Hamiltonian . Koopman operator theory is also used to establish various data-driven analysis for complex dynamics Koopman1 ; EDMD ; Koopman2 . Nonparametric regression techniques for learning interaction kernels KernelLearning1 ; KernelLearning2 ; KernelLearning3 ; KernelLearning4 ; KernelLearning5 ; KernelLearning6 ; Ding2022 are developed for various equations. Flow maps flowMap1 ; flowMap2 ; flowMap3 and kernel flows kernelFlow1 for learning dynamical systems are introduced. Probabilistic/statistical methods, including Bayesian inferences, maximum likelihood methods, Gaussian processes, kernel methods, and Wasserstein distances, are introduced to learn stochastic dynamical systems Felix ; VI4SDE ; LiuSWasserstein ; GP4PDEs1 ; GP4PDEs2 ; Kernel4PDEs . More recently, in order to maintain the physical properties (e.g., invariant quantities for conserved systems or dissipation rates for dissipative systems) of the original system while learning the system, various structure-preserving learning strategies are developed to learn Hamiltonian systems SympNets ; HenonMaps ; MTao ; Hamiltonian1 ; Hamiltonian2 ; Hamiltonian3 ; Hamiltonian4 ; Hamiltonian5 ; SymplecticRNN ; MLforIrreversibleProcess1 , energy dissipative systems dissipativeNN2 ; OnsagerNet ; stat-PINNs , and, more generally, metriplectic systems metriplectic1 ; MLforIrreversibleProcess2 .
@@
However, most existing work establishes the learning framework based on the corresponding governing equations, such as (stochastic) ordinary differential equations (ODEs/SDEs) and partial differential equations (PDEs). The resulting models, however, may fail to preserve fundamental physical principles, such as consistency with thermodynamic laws. Recently, there are growing interesting of learning thermodynamically consistent physical model from variational principles, such as General Equation for Non-Equilibrium Reversible-Irreversible Coupling (GENERIC) formalism hernandez2021structure ; zhang2022gfinns and Onsager principle huang2022variational ; OnsagerNet . These variational principles model complex physical processes by accounting for both energy conservation (in reversible processes) and energy dissipation (in irreversible processes). The key idea behind these variational principle-based learning approaches is to parameterize the physical quantities in the energy-dissipation law using neural networks while constructing loss functions based on equations derived from these principles.
The variational formulation of loss functions has gained increasing attention recently due to its robustness to corrupted or noisy observations. For example, several works have focused on leveraging the weak form of PDEs to construct loss functions for learning the solution of PDEs in forward problems or identifying coefficients in inverse problems DGM ; neuralGalerkin ; weakPINNs ; variationalPINN ; weakSINDy ; weakSINDy4PDE ; weakSINDy4meanfield ; Phase-field_DeepONet . However, most existing work relies on the careful selection of test functions, which can be challenging or even infeasible for high-dimensional problems. Additionally, we note a concurrent independent effort that also aims to address high-dimensional problems by introducing self-test loss functions for learning weak-form operators and gradient flows Lu_self-test . Moreover, the authors in EntropyInformed proposed an entropy-informed learning framework.
The goal of this work is to propose a new learning framework based on the energy-dissipation laws of the target physical systems directly, without relying on the governing equations. Our proposed methods offer several benefits, including robustness to corrupted or noisy observations, straightforward extensions to more general physical systems, and the potential to handle higher-dimensional systems. Moreover, our approach can learn the full dynamics of the system using observations at only three time instances, which not only reduces data requirements but also enables efficient modeling in scenarios where long-time trajectory data are difficult to obtain. While loss functions formulated in the weak form of governing equations are generally more robust to noise than those based on strong formulations, they may struggle to uniquely capture local information and can be challenging to construct for complex systems. Our approach constructs the loss function directly from the energy-dissipation law, enabling effective learning of PDE solutions in forward problems and accurate identification of model parameters in inverse problems, without relying on the explicit form of the governing equations.
@@
In this work, we focus on learning the potential function and noise intensity in one- or two-dimensional generalized diffusions to illustrate our method and explore its performance under different settings. While extending the approach to higher-dimensional problems and other physical systems is straightforward, we leave this direction for future work. The rest of the paper is organized as follows. Section 2 provides a brief introduction to the energetic variational approach for generalized diffusions. In Section 3, we propose a framework for learning the governing laws of the systems based on their energy-dissipation laws, using either continuous data (probability density) or discrete data (SDE particles). Section 4 presents several representative examples to validate the performance of our methods. Finally, we conclude with a brief discussion in Section 5.
@@ 2 Formulation
Before proposing the learning framework, we briefly introduce the energetic variational approach (EnVarA for short) EnVarA1 for generalized diffusions, which plays an important role in our proposed learning frameworks in the next section.
Motivated by non-equilibrium thermodynamics, particularly the seminal work of Rayleigh Rayleigh and Onsager Onsager1 ; Onsager2 , an isothermal and mechanically-closed complex system can be described by an energy-dissipation law
@@
where $E^{{\rm total}}$ is the sum of the kinetic energy $\mathcal{K}$ and the Helmholtz free energy $\mathcal{F}$, and $\Delta$ is the rate of energy dissipation. Based on the energy-dissipation law (1), EnVarA is a unique, well-defined way to derive the dynamics of the underlying system using the least action principle (LAP) and the maximum dissipation principle (MDP). To be more specific, for the Hamiltonian part of the system, one can employ the LAP, taking variation of the action functional $\mathcal{A}({\bm{x}})=\int_{0}^{T}\left(\mathcal{K}-\mathcal{F}\right){\rm d}t$ with respect to $\bm{x}$} (the trajectory in Lagrangian coordinates) EnVarA1 ; arnol2013mathematical , to derive the conservative force, i.e., $\delta\mathcal{A}=\int_{0}^{T}\int_{\Omega}({\rm force}_{\text{iner}}-{\rm force}_{\text{conv}})\cdot\delta{\bm{x}}~{\rm d}{\bm{x}}{\rm d}t.$ Here, $\Omega$ could be a bounded or unbounded domain of $\bm{x}$}, and ${\rm force}_{\text{iner}}$ and ${\rm force}_{\text{conv}}$ are inertial force and conservative force respectively. For the dissipation part, one can apply the MDP, taking the variation of the Onsager dissipation functional $\mathcal{D}$ with respect to the “rate” $\dot{{\bm{x}}}$ ($\dot{{\bm{x}}}$ is the derivative of the trajectory $\bm{x}$} with respect to time $t$), to derive the dissipative force, i.e., $\delta\mathcal{D}=\int_{\Omega}{\rm force}_{\text{diss}}\cdot\delta\dot{{\bm{x}}}~{\rm d}{\bm{x}}$, where the dissipation functional $\mathcal{D}=\frac{1}{2}\triangle$ in the linear response regime Onsager2 and ${\rm force}_{\text{diss}}$ is the dissipative force. Subsequently, the force balance condition connects the conservative force and the dissipation force providing the evolution equation of the studied system
$\frac{\delta\mathcal{D}}{\delta\dot{{\bm{x}}}}=\frac{\delta\mathcal{A}}{\delta\bm{x}}.$ (2)
The EnVarA has been successfully applied to build various mathematical models in physics, chemical engineering, and biology EnVarA_review .
@@
where $W_{t}$ is a standard $n$-dimensional Brownian motion, $\bm{X}_{t}$ and $\bm{a}$ are two $n$-dimensional vectors denoting the state variable at time $t\in\mathbb{R}^{+}\cup\{0\}$ and the drift coefficient respectively, and the noise intensity $\sigma$ is a scalar function. If the stochastic integral of (3) is interpreted as backward Itô integral russo1993forward , one may obtain the following Fokker–Planck equation (See (c) in Remark 2.1):
$f_{t}+\nabla\cdot(\bm{a}f)=\frac{1}{2}\nabla\cdot\left(\sigma^{2}\nabla f\right),$ (4)
@@
According to the fluctuation-dissipation theorem kubo1966fluctuation , the convection coefficient is constrained by
@@
The fluctuation-dissipation theorem ensures the existence of an energy-dissipation law associated with the Fokker–Planck equation (4). It can be shown that the Fokker–Planck equation (4) with the condition (5) satisfies energy-dissipation law:
$\frac{d\mathcal{F}[f]}{dt}=-\int_{\Omega}\frac{f}{\sigma^{2}/2}|\bm{u}|^{2}d\bm{x},$ (6)
@@
Here, $\bm{u}$} is a certain average velocity of all stochastic trajectories and $\mathcal{F}[f]$ is the free energy given by
${\bm{u}}=-\frac{\sigma^{2}}{2}\nabla(\ln f+\psi),\quad\mathcal{F}[f]:=\int_{\Omega}\left[f\ln f+\psi f\right]d\bm{x}\ .$ (8)
From a modeling perspective, one can derive the evolution equation (4) from the energy-dissipation law (6) by the general framework of EnVarA EnVarA1 . Note that
$\mathcal{K}=0,\qquad\mathcal{F}=\int_{\Omega}\left[f\ln f+\psi f\right]d\bm{x},\qquad\mathcal{D}=\frac{1}{2}\int_{\Omega}\frac{f}{\sigma^{2}/2}|\bm{u}|^{2}d\bm{x}.$ (9)
@@
which is the velocity derived from the energy-dissipation law (6). Combining with the continuity equation (7), one can obtain the Fokker–Planck equation (4) with $\bm{a}$} given by (5). An advantage of deriving the governing equation from an energy-dissipation law is that the resulting system is automatically thermodynamically consistent, meaning it satisfies the fluctuation-dissipation theorem in this case. We refer the interested reader to EnVarA1 ; hu2024energetic and the references therein for more details.
@@
Different interpretations of the stochastic integral of (3) leads to different energy-dissipation law EnVarA1 . To be more specific, writing a Taylor expansion of probability distribution function $f(\bm{x},t)$, one may obtain the following PDEs EnVarA1 :